# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Skill: read-questions-with-short-vowel-words
Description: Answer the question.
Can a wig jog?
Choices: [(a) yes, (b) no]
Answer index: 1
Subject: Language Arts
@@
Skill: choose-the-uppercase-letter-that-matches-review
Description: Pick the uppercase letter that matches.
x
Choices: [(a) X, (b) W]
Answer index: 0
Subject: Language Arts
Skill: find-the-action-verb
Description: Find the verb. A verb shows an action.
Choices: [(a) tree, (b) grow]
Answer index: 1
Subject: Language Arts
@@
Skill: choose-the-lowercase-letter-that-matches-review
Description: Pick the lowercase letter that matches.
V
Choices: [(a) n, (b) v]
Answer index: 1
Subject: Language Arts
Skill: find-the-complete-sentence
Description: Find the complete sentence.
Choices: [(a) car, (b) We ate the cake.]
Answer index: 1
@@
Subject: Language Arts
Skill: which-word-is-not-like-the-others
Description: Which word is not like the others?
Choices: [(a) big, (b) tall, (c) little, (d) nice]
Answer index: 3
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-sight-word-sets-1-2-3
Description: Complete the sentence.
I did not want to trick ___.
Choices: [(a) has, (b) as, (c) them]
Answer index: 2
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7
Description: Complete the sentence.
I ate a hot dog ___ went to bed.
Choices: [(a) four, (b) more, (c) then]
Answer index: 2
Subject: Language Arts
Skill: use-the-correct-modal-verb
Description: Select the modal verb that best completes the sentence.
The coach ___ let Mark join the team if it weren’t so late in the season.
Choices: [(a) would, (b) will]
Answer index: 0
Subject: Language Arts
Skill: use-the-perfect-verb-tenses
Description: Complete the sentence with the correct helping verb or verbs.
When I walked by Sophia’s desk, I saw that she ___ fallen asleep!
Choices: [(a) had, (b) will have]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meaning-of-idioms-from-context-set-2
Description: What is the meaning of **get one’s feet wet**?
Mrs. Tyler is a professional singer who **got her feet wet** by singing in the school choir as a kid.
Choices: [(a) to get soaked or messy, (b) to start learning something]
Answer index: 1
Subject: Language Arts
Skill: choose-between-the-past-tense-and-past-participle
Description: Complete the sentence with the correct form of the verb.
I swept up crumbs from the cookie Barry had ___.
Choices: [(a) eaten, (b) ate]
Answer index: 0
Subject: Language Arts
Skill: identify-the-meaning-of-idioms-and-adages-set-2
Description: What is the meaning of **if you can’t stand the heat, get out of the kitchen**?
Choices: [(a) follow the customs of the place where you are, (b) don’t take on a job or task if it’s too much for you]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-declarative-interrogative-imperative-or-exclamatory
Description: What kind of sentence is this?
Nora is reading a magazine about skateboarding.
Choices: [(a) interrogative, (b) declarative]
Answer index: 1
Subject: Language Arts
Skill: use-the-progressive-verb-tenses
Description: Complete the sentence with the correct helping verb or verbs.
"We ___ eating dinner at seven o’clock, so don’t be late," Mrs. Christensen said.
Choices: [(a) were, (b) will be]
Answer index: 1
Subject: Language Arts
Skill: what-does-the-modal-verb-show
Description: What does the modal verb show? Select the best answer choice.
Bernie and Gary **will** visit Yosemite National Park next summer.
Choices: [(a) a request for permission, (b) future action]
Answer index: 1
Subject: Language Arts
Skill: identify-the-meaning-of-idioms-and-adages-set-1
Description: What is the meaning of **cut corners**?
Choices: [(a) to accidentally take the edges off something, (b) to take shortcuts when working]
Answer index: 1
Subject: Language Arts
Skill: use-thesaurus-entries
Description: Look at this thesaurus entry:
*abstain**
Synonyms: deny, forgo, give up, quit, refrain
Which is something you **abstain** from?
Choices: [(a) something you almost always do, (b) something you choose not to do]
Answer index: 1
Subject: Language Arts
Skill: words-with-sub
Description: Which might be a **subtopic** within a discussion of the Revolutionary War?
Choices: [(a) the Declaration of Independence, (b) U.S. history]
Answer index: 0
Subject: Language Arts
Skill: words-with-pre
Description: What helps **prevent** a cold?
Choices: [(a) washing your hands frequently, (b) taking cold medicine]
Answer index: 0
Subject: Language Arts
Skill: words-with-less
Description: Why might you make a **careless** mistake in your work?
Choices: [(a) because you aren’t paying attention, (b) because you don’t understand something]
Answer index: 0
Subject: Language Arts
Skill: words-with-un-dis-in-im-and-non
Description: Which is an **infrequent** visitor?
Choices: [(a) one you often see, (b) one you rarely see]
Answer index: 1
Subject: Language Arts
Skill: words-with-re
Description: Why might you **reattempt** a challenge?
Choices: [(a) because you didn’t succeed the first time, (b) because it’s a new one]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-pair-of-correlative-conjunctions
Description: Complete the sentence with the correct correlative conjunctions.
Kurt’s horse is in ___ the barn ___ the pasture right now.
Choices: [(a) not only . . . but also, (b) either . . . or]
Answer index: 1
Subject: Language Arts
Skill: words-with-ful
Description: How would you feel after a **restful** day?
Choices: [(a) recharged, (b) exhausted]
Answer index: 0
Subject: Language Arts
Skill: compound-subjects-and-objects-with-personal-pronouns
Description: Select the correct pronoun to complete the sentence.
At the garage sale, ___ and Warren hoped to find some good books.
Choices: [(a) she, (b) her]
Answer index: 0
Subject: Language Arts
Skill: determine-the-themes-of-short-stories
Description: Read the following story.
Paolo walked out on the dock and looked down into the water. "I only like to swim in pools. I’m not swimming in there," he said.
"Well, I’m going to," his mom said. "There’s not much else to do here!"
Paolo’s little sister, Celia, was standing next to him. "If Paolo’s not going in, I’m not swimming in there either," Celia said.
Paolo’s mom looked at Paolo with a raised eyebrow. "That’s fine," she said. "You two don’t have to swim. But remember, there’s no internet here and no TV."
Paolo marched back to the cabin and sat in a lawn chair. He crossed his arms and frowned. "This is going to be an awful vacation", he thought. No Wi-Fi, nothing to do.
Celia followed him, sitting in a chair and crossing her arms, too. "I’m bored!" Celia said.
Meanwhile, their mom had gone into the cabin to change. Soon she emerged, carrying an armful of goggles and towels. "I’m going for a dip. Are you "sure" you don’t want to go?" she asked.
Celia glanced at Paolo.
"I’m sure," said Paolo.
"No way!" added Celia.
Paolo sat in the chair and stared across the lake. Celia did too, but soon she grew restless. She got up, walked up and down the dock, and finally sat down, watching her mom.
Seeing this, Paolo felt bad. He went inside and put on his bathing suit. Then he walked out to the dock where Celia was sitting. When their mom swam up to them, Paolo apologized. "I’ll swim with you, Mom. I’m sorry I was crabby. The water actually looks nice," he said.
Celia looked at Paolo. Then she ran to the cabin to change into her swimsuit, too.
Which of the following best describes the main theme or lesson of the story?
Choices: [(a) Set a good example for people who look up to you., (b) People today take technology for granted.]
Answer index: 0
@@
Subject: Language Arts
Skill: words-with-able-and-ible
Description: Which item is not a **collectible**?
Choices: [(a) an ordinary spiral notebook, (b) a rare comic book]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-frequently-confused-word
Description: Complete the text with the correct word.
@@
Rhianna made a wire sculpture of a tree for her art project. She used steel wool to make ___ leaves.
Choices: [(a) its, (b) it’s]
Answer index: 0
Subject: Language Arts
Skill: words-with-mis
Description: Which type of behavior shouldn’t cause **mistrust**?
Choices: [(a) suspicious behavior, (b) honest behavior]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-is-more-formal
Description: Which sentence is more formal?
@@
Choices: [(a) Visitors are advised that admission to the aircraft museum has been raised by two dollars., (b) FYI, admission to the aircraft museum has been raised by two dollars.]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-simple-compound-or-complex
Description: Which is a **simple sentence**?
@@
Choices: [(a) Although I would rather visit Italy, I’m excited about going to Panama this summer., (b) The thermometer and the cough syrup are in the medicine cabinet next to the cotton balls.]
Answer index: 1
Subject: Language Arts
Skill: choose-topic-sentences-for-expository-paragraphs
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ In fact, more than half of the planet is covered just by oceans. Earth also has many lakes, rivers, and swamps full of freshwater. But there’s even more freshwater that is frozen in glaciers and icecaps. Add to that all of the water that’s underground or in the air, and you’ve got one wet planet!
Choices: [(a) Most of the Earth is covered in water., (b) The highest point on Earth is Mount Everest.]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-a-statement-question-command-or-exclamation
Description: What kind of sentence is this?
Bobby gave me the best present!
Choices: [(a) an exclamation, (b) a statement]
Answer index: 0
Subject: Language Arts
Skill: does-the-adjective-tell-you-what-kind-or-how-many
@@
Description: Look at the adjective in **bold**. Does it tell you "what kind" or "how many"?
Maria will need **seven** stamps.
Choices: [(a) what kind, (b) how many]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-similes
Description: What is the meaning of the simile?
Just as a soft breeze rocks the trees, Mrs. Middleton rocked the baby to sleep.
Choices: [(a) Mrs. Middleton rocked the baby outside., (b) Mrs. Middleton rocked the baby gently.]
Answer index: 1
Subject: Language Arts
Skill: is-the-noun-singular-or-plural
Description: Is the noun in bold singular or plural?
Lamar shouted when his **foot** touched the hot water.
Choices: [(a) singular, (b) plural]
Answer index: 0
Subject: Language Arts
Skill: use-the-correct-article-a-or-an
Description: Complete the sentence with the best article.
Mrs. Casey keeps the tools in ___ shed.
Choices: [(a) an, (b) a]
Answer index: 1
Subject: Language Arts
Skill: determine-the-themes-of-myths-fables-and-folktales
Description: Read the following fable by Aesop.
A fox one day spotted a beautiful bunch of ripe grapes hanging from a vine. The grapes seemed ready to burst with juice. The fox’s mouth watered, and he gazed longingly at them.
The bunch hung up high, far out of the fox’s reach. He tried to jump and grab it. The first time he jumped, he was still far away. So he walked off a short distance and took a running leap. But he fell short once more. Again and again he tried, but he simply could not reach the grapes.
After a long time, he sat down and looked at the grapes in anger.
"What a fool I am," he said. "I am wearing myself out, and for what? This is just a bunch of terrible, sour grapes."
And away he walked with his nose in the air.
What is the main theme or lesson of the fable?
Choices: [(a) Sometimes fruit is sour, not sweet., (b) People pretend they don’t want things they can’t have.]
Answer index: 1
Subject: Language Arts
Skill: to-be-use-the-correct-form
@@
Description: Complete the sentence with the correct **present tense** form of **to be**.
They ___ upstairs in Sofia’s bedroom.
Choices: [(a) are, (b) am]
Answer index: 0
Subject: Language Arts
Skill: read-about-famous-places
Description: Review the third paragraph.
The Golden Gate Bridge is one of the most famous bridges in the world. Its tall frame, two large towers, and orange color make the bridge easy to recognize. Yet as the Golden Gate Bridge was being built, people disagreed about what the bridge should look like.
People were worried that San Francisco’s thick fog would make the bridge hard to see. The U.S. Navy wanted the bridge to stand out so ships and airplanes wouldn’t collide with it. They wanted the bridge to have wide black and yellow stripes, like a bumblebee! Irving Morrow was one of the bridge designers. He did not like the idea of a striped bridge. Morrow did not want a plain black or gray bridge either, though. He thought a dark bridge would make the Golden Gate seem smaller than it was.
When the bridge’s parts were delivered to San Francisco, they had an orange color. The bridge workers were supposed to paint a new color over the orange. But Morrow was still unsure of what color that would be. He looked at the orange bridge parts and decided that the color would be perfect! Morrow thought the orange would stand out well against the ocean and the sky.
Finally, the difficult decision had been made. A red-orange Golden Gate Bridge was bright enough for ships and planes to see. The color made the bridge appear grand and tall. Over the years, fresh paint has been needed. But the color remains the same orange that was supposed to be painted over many years ago.
Based on the text, why did Irving Morrow think orange would be a good color for the Golden Gate Bridge?
Choices: [(a) The parts were already orange, so no extra work would be needed., (b) The color would make the bridge stand out from the sky and ocean., (c) Using orange paint on orange parts would be cheaper.]
Answer index: 1
Subject: Language Arts
Skill: does-the-adverb-tell-you-how-when-or-where
Description: Look at the adverb in bold. Does it tell you "how" or "where"?
Mr. Olson lives **upstairs**.
Choices: [(a) how, (b) where]
Answer index: 1
Subject: Language Arts
Skill: which-word-is-a-noun
Description: Is the word in bold a noun?
The captain turned the ship around and **sailed** north.
Choices: [(a) yes, (b) no]
Answer index: 1
@@
Subject: Language Arts
Skill: identify-the-irregular-past-tense-ii
Description: Which sentence is in the past tense?
Choices: [(a) Susan **feeds** her bird., (b) Susan **fed** her bird.]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-is-in-the-regular-past-tense
Description: Which sentence is in the past tense?
Choices: [(a) Cassie **acted** happy., (b) We **discover** hidden treasures.]
Answer index: 0
Subject: Language Arts
Skill: which-sentence-uses-an-antonym
@@
Description: Read the sentence. Select the sentence that uses an antonym in place of the word in bold.
Bella’s dad has always been **short**.
Choices: [(a) Bella’s dad has always been **tall**., (b) Bella’s dad has always been **shy**.]
Answer index: 0
Subject: Language Arts
Skill: use-action-verbs
Description: Complete the sentence with the action verb.
Mrs. Carlson ___ nuts into the cookie batter.
Choices: [(a) mixes, (b) butter]
Answer index: 0
Subject: Language Arts
Skill: pronoun-verb-agreement
Description: Complete the sentence with the best **verb**.
He ___ on television.
Choices: [(a) appears, (b) appear]
Answer index: 0
Subject: Language Arts
Skill: make-predictions-about-a-story
Description: Read this passage from a story.
Anna and Andrew had just moved in to their new house. "Let’s explore," Anna said. They opened all the closets in the bedrooms. Then they found the door to the attic. Their dad noticed them opening the door. "Go ahead," he said. "It’s full of old junk."
In the attic, Anna and Andrew saw an old dresser and some boards. Then they noticed a very large metal box. They walked over to look at it more closely.
Based on the passage, which is more likely to happen next?
Choices: [(a) Anna and Andrew open the metal box., (b) Anna and Andrew open the dresser drawers.]
Answer index: 0
Subject: Language Arts
Skill: greetings-and-closings-of-letters
Description: Which closing is correct for a letter?
Choices: [(a) Thanks, Jim, (b) thanks, Jim]
Answer index: 0
Subject: Language Arts
Skill: identify-base-words-prefixes-and-suffixes
Description: What is the base word in **dampness**?
Choices: [(a) damp, (b) ness]
Answer index: 0
Subject: Language Arts
Skill: which-sentence-has-the-same-meaning
Description: Which sentence has the same meaning?
Mrs. Duncan **closed** her bedroom door and went to bed.
Choices: [(a) Mrs. Duncan **shut** her bedroom door and went to bed., (b) Mrs. Duncan **painted** her bedroom door and went to bed.]
Answer index: 0
Subject: Language Arts
Skill: identify-the-authors-purpose-passages
Description: Read the text below.
Vote for Bob Mueller for city council. He’ll stand up for the needs of our town’s hardworking citizens.
What is the text’s most likely purpose?
Choices: [(a) to persuade, (b) to inform, (c) to entertain]
Answer index: 0
Subject: Language Arts
Skill: choose-topic-sentences-for-narrative-paragraphs
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ First, my parents took me to the store to buy some seeds. Then, we took the seeds home, and I planted them in my backyard. I watered them every day, and soon the plants started to grow. After several weeks, I started to see little vegetables pop up. By the end of the summer, I had a garden full of green peppers, beans, corn, and carrots.
Choices: [(a) Last spring, I planted an amazing vegetable garden., (b) Fresh vegetables are my favorite topping to put on pizza.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-text-that-matches-the-writers-purpose
Description: You want to **inform** someone **about blueberries**. "Inform" means "teach someone about something". What should you write?

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: read-questions-with-short-vowel-words Description: Answer the question. Can a wig jog? Choices: [(a) yes, (b) no] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-uppercase-letter-that-matches-review Description: Pick the uppercase letter that matches. x Choices: [(a) X, (b) W] Answer index: 0
Subject: Language Arts Skill: find-the-action-verb Description: Find the verb. A verb shows an action. Choices: [(a) tree, (b) grow] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-lowercase-letter-that-matches-review Description: Pick the lowercase letter that matches. V Choices: [(a) n, (b) v] Answer index: 1
Subject: Language Arts Skill: find-the-complete-sentence Description: Find the complete sentence. Choices: [(a) car, (b) We ate the cake.] Answer index: 1
@@
Subject: Language Arts Skill: which-word-is-not-like-the-others Description: Which word is not like the others? Choices: [(a) big, (b) tall, (c) little, (d) nice] Answer index: 3
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word-sets-1-2-3 Description: Complete the sentence. I did not want to trick ___. Choices: [(a) has, (b) as, (c) them] Answer index: 2
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7 Description: Complete the sentence. I ate a hot dog ___ went to bed. Choices: [(a) four, (b) more, (c) then] Answer index: 2
Subject: Language Arts Skill: use-the-correct-modal-verb Description: Select the modal verb that best completes the sentence. The coach ___ let Mark join the team if it weren’t so late in the season. Choices: [(a) would, (b) will] Answer index: 0
Subject: Language Arts Skill: use-the-perfect-verb-tenses Description: Complete the sentence with the correct helping verb or verbs. When I walked by Sophia’s desk, I saw that she ___ fallen asleep! Choices: [(a) had, (b) will have] Answer index: 0
Subject: Language Arts Skill: determine-the-meaning-of-idioms-from-context-set-2 Description: What is the meaning of **get one’s feet wet**? Mrs. Tyler is a professional singer who **got her feet wet** by singing in the school choir as a kid. Choices: [(a) to get soaked or messy, (b) to start learning something] Answer index: 1
Subject: Language Arts Skill: choose-between-the-past-tense-and-past-participle Description: Complete the sentence with the correct form of the verb. I swept up crumbs from the cookie Barry had ___. Choices: [(a) eaten, (b) ate] Answer index: 0
Subject: Language Arts Skill: identify-the-meaning-of-idioms-and-adages-set-2 Description: What is the meaning of **if you can’t stand the heat, get out of the kitchen**? Choices: [(a) follow the customs of the place where you are, (b) don’t take on a job or task if it’s too much for you] Answer index: 1
Subject: Language Arts Skill: is-the-sentence-declarative-interrogative-imperative-or-exclamatory Description: What kind of sentence is this? Nora is reading a magazine about skateboarding. Choices: [(a) interrogative, (b) declarative] Answer index: 1
Subject: Language Arts Skill: use-the-progressive-verb-tenses Description: Complete the sentence with the correct helping verb or verbs. "We ___ eating dinner at seven o’clock, so don’t be late," Mrs. Christensen said. Choices: [(a) were, (b) will be] Answer index: 1
Subject: Language Arts Skill: what-does-the-modal-verb-show Description: What does the modal verb show? Select the best answer choice. Bernie and Gary **will** visit Yosemite National Park next summer. Choices: [(a) a request for permission, (b) future action] Answer index: 1
Subject: Language Arts Skill: identify-the-meaning-of-idioms-and-adages-set-1 Description: What is the meaning of **cut corners**? Choices: [(a) to accidentally take the edges off something, (b) to take shortcuts when working] Answer index: 1
Subject: Language Arts Skill: use-thesaurus-entries Description: Look at this thesaurus entry: **abstain** Synonyms: deny, forgo, give up, quit, refrain Which is something you **abstain** from? Choices: [(a) something you almost always do, (b) something you choose not to do] Answer index: 1
Subject: Language Arts Skill: words-with-sub Description: Which might be a **subtopic** within a discussion of the Revolutionary War? Choices: [(a) the Declaration of Independence, (b) U.S. history] Answer index: 0
Subject: Language Arts Skill: words-with-pre Description: What helps **prevent** a cold? Choices: [(a) washing your hands frequently, (b) taking cold medicine] Answer index: 0
Subject: Language Arts Skill: words-with-less Description: Why might you make a **careless** mistake in your work? Choices: [(a) because you aren’t paying attention, (b) because you don’t understand something] Answer index: 0
Subject: Language Arts Skill: words-with-un-dis-in-im-and-non Description: Which is an **infrequent** visitor? Choices: [(a) one you often see, (b) one you rarely see] Answer index: 1
Subject: Language Arts Skill: words-with-re Description: Why might you **reattempt** a challenge? Choices: [(a) because you didn’t succeed the first time, (b) because it’s a new one] Answer index: 0
Subject: Language Arts Skill: use-the-correct-pair-of-correlative-conjunctions Description: Complete the sentence with the correct correlative conjunctions. Kurt’s horse is in ___ the barn ___ the pasture right now. Choices: [(a) not only . . . but also, (b) either . . . or] Answer index: 1
Subject: Language Arts Skill: words-with-ful Description: How would you feel after a **restful** day? Choices: [(a) recharged, (b) exhausted] Answer index: 0
Subject: Language Arts Skill: compound-subjects-and-objects-with-personal-pronouns Description: Select the correct pronoun to complete the sentence. At the garage sale, ___ and Warren hoped to find some good books. Choices: [(a) she, (b) her] Answer index: 0
Subject: Language Arts Skill: determine-the-themes-of-short-stories Description: Read the following story. Paolo walked out on the dock and looked down into the water. "I only like to swim in pools. I’m not swimming in there," he said. "Well, I’m going to," his mom said. "There’s not much else to do here!" Paolo’s little sister, Celia, was standing next to him. "If Paolo’s not going in, I’m not swimming in there either," Celia said. Paolo’s mom looked at Paolo with a raised eyebrow. "That’s fine," she said. "You two don’t have to swim. But remember, there’s no internet here and no TV." Paolo marched back to the cabin and sat in a lawn chair. He crossed his arms and frowned. "This is going to be an awful vacation", he thought. No Wi-Fi, nothing to do. Celia followed him, sitting in a chair and crossing her arms, too. "I’m bored!" Celia said. Meanwhile, their mom had gone into the cabin to change. Soon she emerged, carrying an armful of goggles and towels. "I’m going for a dip. Are you "sure" you don’t want to go?" she asked. Celia glanced at Paolo. "I’m sure," said Paolo. "No way!" added Celia. Paolo sat in the chair and stared across the lake. Celia did too, but soon she grew restless. She got up, walked up and down the dock, and finally sat down, watching her mom. Seeing this, Paolo felt bad. He went inside and put on his bathing suit. Then he walked out to the dock where Celia was sitting. When their mom swam up to them, Paolo apologized. "I’ll swim with you, Mom. I’m sorry I was crabby. The water actually looks nice," he said. Celia looked at Paolo. Then she ran to the cabin to change into her swimsuit, too. Which of the following best describes the main theme or lesson of the story? Choices: [(a) Set a good example for people who look up to you., (b) People today take technology for granted.] Answer index: 0
@@
Subject: Language Arts Skill: words-with-able-and-ible Description: Which item is not a **collectible**? Choices: [(a) an ordinary spiral notebook, (b) a rare comic book] Answer index: 0
@@
Subject: Language Arts Skill: use-the-correct-frequently-confused-word Description: Complete the text with the correct word. Rhianna made a wire sculpture of a tree for her art project. She used steel wool to make ___ leaves. Choices: [(a) its, (b) it’s] Answer index: 0
Subject: Language Arts Skill: words-with-mis Description: Which type of behavior shouldn’t cause **mistrust**? Choices: [(a) suspicious behavior, (b) honest behavior] Answer index: 1
@@
Subject: Language Arts Skill: which-sentence-is-more-formal Description: Which sentence is more formal? Choices: [(a) Visitors are advised that admission to the aircraft museum has been raised by two dollars., (b) FYI, admission to the aircraft museum has been raised by two dollars.] Answer index: 0
@@
Subject: Language Arts Skill: is-the-sentence-simple-compound-or-complex Description: Which is a **simple sentence**? Choices: [(a) Although I would rather visit Italy, I’m excited about going to Panama this summer., (b) The thermometer and the cough syrup are in the medicine cabinet next to the cotton balls.] Answer index: 1
@@
Subject: Language Arts Skill: choose-topic-sentences-for-expository-paragraphs Description: Read the paragraph. Then, add the better topic sentence. ___ In fact, more than half of the planet is covered just by oceans. Earth also has many lakes, rivers, and swamps full of freshwater. But there’s even more freshwater that is frozen in glaciers and icecaps. Add to that all of the water that’s underground or in the air, and you’ve got one wet planet! Choices: [(a) Most of the Earth is covered in water., (b) The highest point on Earth is Mount Everest.] Answer index: 0
Subject: Language Arts Skill: is-the-sentence-a-statement-question-command-or-exclamation Description: What kind of sentence is this? Bobby gave me the best present! Choices: [(a) an exclamation, (b) a statement] Answer index: 0
@@
Subject: Language Arts Skill: does-the-adjective-tell-you-what-kind-or-how-many Description: Look at the adjective in **bold**. Does it tell you "what kind" or "how many"? Maria will need **seven** stamps. Choices: [(a) what kind, (b) how many] Answer index: 1
Subject: Language Arts Skill: determine-the-meanings-of-similes Description: What is the meaning of the simile? Just as a soft breeze rocks the trees, Mrs. Middleton rocked the baby to sleep. Choices: [(a) Mrs. Middleton rocked the baby outside., (b) Mrs. Middleton rocked the baby gently.] Answer index: 1
Subject: Language Arts Skill: is-the-noun-singular-or-plural Description: Is the noun in bold singular or plural? Lamar shouted when his **foot** touched the hot water. Choices: [(a) singular, (b) plural] Answer index: 0
Subject: Language Arts Skill: use-the-correct-article-a-or-an Description: Complete the sentence with the best article. Mrs. Casey keeps the tools in ___ shed. Choices: [(a) an, (b) a] Answer index: 1
Subject: Language Arts Skill: determine-the-themes-of-myths-fables-and-folktales Description: Read the following fable by Aesop. A fox one day spotted a beautiful bunch of ripe grapes hanging from a vine. The grapes seemed ready to burst with juice. The fox’s mouth watered, and he gazed longingly at them. The bunch hung up high, far out of the fox’s reach. He tried to jump and grab it. The first time he jumped, he was still far away. So he walked off a short distance and took a running leap. But he fell short once more. Again and again he tried, but he simply could not reach the grapes. After a long time, he sat down and looked at the grapes in anger. "What a fool I am," he said. "I am wearing myself out, and for what? This is just a bunch of terrible, sour grapes." And away he walked with his nose in the air. What is the main theme or lesson of the fable? Choices: [(a) Sometimes fruit is sour, not sweet., (b) People pretend they don’t want things they can’t have.] Answer index: 1
@@
Subject: Language Arts Skill: to-be-use-the-correct-form Description: Complete the sentence with the correct **present tense** form of **to be**. They ___ upstairs in Sofia’s bedroom. Choices: [(a) are, (b) am] Answer index: 0
Subject: Language Arts Skill: read-about-famous-places Description: Review the third paragraph. The Golden Gate Bridge is one of the most famous bridges in the world. Its tall frame, two large towers, and orange color make the bridge easy to recognize. Yet as the Golden Gate Bridge was being built, people disagreed about what the bridge should look like. People were worried that San Francisco’s thick fog would make the bridge hard to see. The U.S. Navy wanted the bridge to stand out so ships and airplanes wouldn’t collide with it. They wanted the bridge to have wide black and yellow stripes, like a bumblebee! Irving Morrow was one of the bridge designers. He did not like the idea of a striped bridge. Morrow did not want a plain black or gray bridge either, though. He thought a dark bridge would make the Golden Gate seem smaller than it was. When the bridge’s parts were delivered to San Francisco, they had an orange color. The bridge workers were supposed to paint a new color over the orange. But Morrow was still unsure of what color that would be. He looked at the orange bridge parts and decided that the color would be perfect! Morrow thought the orange would stand out well against the ocean and the sky. Finally, the difficult decision had been made. A red-orange Golden Gate Bridge was bright enough for ships and planes to see. The color made the bridge appear grand and tall. Over the years, fresh paint has been needed. But the color remains the same orange that was supposed to be painted over many years ago. Based on the text, why did Irving Morrow think orange would be a good color for the Golden Gate Bridge? Choices: [(a) The parts were already orange, so no extra work would be needed., (b) The color would make the bridge stand out from the sky and ocean., (c) Using orange paint on orange parts would be cheaper.] Answer index: 1
Subject: Language Arts Skill: does-the-adverb-tell-you-how-when-or-where Description: Look at the adverb in bold. Does it tell you "how" or "where"? Mr. Olson lives **upstairs**. Choices: [(a) how, (b) where] Answer index: 1
Subject: Language Arts Skill: which-word-is-a-noun Description: Is the word in bold a noun? The captain turned the ship around and **sailed** north. Choices: [(a) yes, (b) no] Answer index: 1
@@
Subject: Language Arts Skill: identify-the-irregular-past-tense-ii Description: Which sentence is in the past tense? Choices: [(a) Susan **feeds** her bird., (b) Susan **fed** her bird.] Answer index: 1
Subject: Language Arts Skill: which-sentence-is-in-the-regular-past-tense Description: Which sentence is in the past tense? Choices: [(a) Cassie **acted** happy., (b) We **discover** hidden treasures.] Answer index: 0
@@
Subject: Language Arts Skill: which-sentence-uses-an-antonym Description: Read the sentence. Select the sentence that uses an antonym in place of the word in bold. Bella’s dad has always been **short**. Choices: [(a) Bella’s dad has always been **tall**., (b) Bella’s dad has always been **shy**.] Answer index: 0
Subject: Language Arts Skill: use-action-verbs Description: Complete the sentence with the action verb. Mrs. Carlson ___ nuts into the cookie batter. Choices: [(a) mixes, (b) butter] Answer index: 0
Subject: Language Arts Skill: pronoun-verb-agreement Description: Complete the sentence with the best **verb**. He ___ on television. Choices: [(a) appears, (b) appear] Answer index: 0
Subject: Language Arts Skill: make-predictions-about-a-story Description: Read this passage from a story. Anna and Andrew had just moved in to their new house. "Let’s explore," Anna said. They opened all the closets in the bedrooms. Then they found the door to the attic. Their dad noticed them opening the door. "Go ahead," he said. "It’s full of old junk." In the attic, Anna and Andrew saw an old dresser and some boards. Then they noticed a very large metal box. They walked over to look at it more closely. Based on the passage, which is more likely to happen next? Choices: [(a) Anna and Andrew open the metal box., (b) Anna and Andrew open the dresser drawers.] Answer index: 0
Subject: Language Arts Skill: greetings-and-closings-of-letters Description: Which closing is correct for a letter? Choices: [(a) Thanks, Jim, (b) thanks, Jim] Answer index: 0
Subject: Language Arts Skill: identify-base-words-prefixes-and-suffixes Description: What is the base word in **dampness**? Choices: [(a) damp, (b) ness] Answer index: 0
Subject: Language Arts Skill: which-sentence-has-the-same-meaning Description: Which sentence has the same meaning? Mrs. Duncan **closed** her bedroom door and went to bed. Choices: [(a) Mrs. Duncan **shut** her bedroom door and went to bed., (b) Mrs. Duncan **painted** her bedroom door and went to bed.] Answer index: 0
Subject: Language Arts Skill: identify-the-authors-purpose-passages Description: Read the text below. Vote for Bob Mueller for city council. He’ll stand up for the needs of our town’s hardworking citizens. What is the text’s most likely purpose? Choices: [(a) to persuade, (b) to inform, (c) to entertain] Answer index: 0
@@
Subject: Language Arts Skill: choose-topic-sentences-for-narrative-paragraphs Description: Read the paragraph. Then, add the better topic sentence. ___ First, my parents took me to the store to buy some seeds. Then, we took the seeds home, and I planted them in my backyard. I watered them every day, and soon the plants started to grow. After several weeks, I started to see little vegetables pop up. By the end of the summer, I had a garden full of green peppers, beans, corn, and carrots. Choices: [(a) Last spring, I planted an amazing vegetable garden., (b) Fresh vegetables are my favorite topping to put on pizza.] Answer index: 0